### v2

# Large Language Models Reflect Human Citation Patterns with a Heightened Citation Bias

Abstract
Citation practices are crucial in shaping the structure of scientific knowledge, yet they are often influenced by contemporary norms and biases. The emergence of Large Language Models (LLMs) like GPT-4 introduces a new dynamic to these practices. Interestingly, the characteristics and potential biases of references recommended by LLMs that entirely rely on their parametric knowledge, and not on search or retrieval-augmented generation, remain unexplored. Here, we analyze these characteristics in an experiment using a dataset of 166 papers from AAAI, NeurIPS, ICML, and ICLR, published after GPT-4’s knowledge cut-off date, encompassing 3,066 references in total. In our experiment, GPT-4 was tasked with suggesting scholarly references for the anonymized in-text citations within these papers. Our findings reveal a remarkable similarity between human and LLM citation patterns, but with a more pronounced high citation bias in GPT-4, which persists even after controlling for publication year, title length, number of authors, and venue. Additionally, we observe a large consistency between the characteristics of GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. By analyzing citation graphs, we show that the references recommended by GPT-4 are embedded in the relevant citation context, suggesting an even deeper conceptual internalization of the citation networks. While LLMs can aid in citation generation, they may also amplify existing biases and introduce new ones, potentially skewing scientific knowledge dissemination. Our results underscore the need for identifying the model’s biases and for developing balanced methods to interact with LLMs in general. [^0] Data and code are available at https://github.com/AndresAlgaba/LLM_citation_patterns.
1 Introduction
Figure 1: Overview of our experiment evaluating the characteristics and biases of the references generated by GPT-4, when tasked to suggest references for anonymized in-text citations. We collect $166$ papers from the cs.LG category on arXiv which are published in the main tracks of AAAI, NeurIPS, ICML, and ICLR, and only appeared available online after GPT-4-0613’s knowledge cut-off date. We split the main content, which includes the author information, conference information, abstract, and introduction, from the ground truth references. GPT-4 is prompted to generate suggestions of scholarly references for the anonymized in-text citations in the main content. We verify the existence of the generated references via Semantic Scholar and compare the characteristics, such as title length, publication year, venue, and number of authors, of the existing and non-existent generated references with the ground truth. For the existing generated references, we also compare additional characteristics, such as the number of citations and references, and analyze the properties of their citation networks.
Large Language Models (LLMs) have revolutionized natural language understanding and generation, driving scientific research forward by assisting in all steps of the scientific process, ranging from identifying research gaps to accelerating complex data analysis [3, 10, 24, 28, 40, 44]. One particularly interesting application is the generation of suggestions for appropriate scholarly references [27, 35]. Yet, without the aid of web browsing or retrieval-augmented generation, these models rely entirely on their parametric knowledge encapsulated during their (pre-)training [5, 6, 14, 38, 41]. Our research focuses on this intrinsic citation behavior of GPT-4, exploring how the model recommends references based on its training data, and highlighting the potential biases that arise from this internalized knowledge [1, 23].
Biases in citation practices have long been a subject of scrutiny in the scientific community [8, 29]. Common biases include a preference for recent publications [4], shorter titles [20], and high-profile publication venues [18]. Moreover, a well-documented phenomenon is the “Matthew effect,” where highly cited papers tend to accumulate even more citations [36]. The number of authors on a paper can also influence its likelihood of being cited, with solo and small author groups often cited less frequently than larger teams [9]. By examining how these biases manifest in the GPT-4 generated references, we aim to uncover the underlying patterns and their potential to amplify existing biases and introduce new ones, potentially skewing scientific knowledge dissemination.
In our experiment, we let GPT-4 suggest scholarly references for anonymized in-text citations within a paper and compare the characteristics and citation networks of the GPT-4 generated references against the ground truth. We provide a comprehensive analysis of $166$ papers which are published in the main tracks of AAAI, NeurIPS, ICML, and ICLR and only available online after GPT-4-0613’s knowledge cut-off date. We find that GPT-4 exhibits strong preferences for highly cited papers, which persists even after controlling for multiple confounding factors such as publication year, title length, venue, and number of authors. Additionally, we observe a large consistency between GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. By analyzing citation graphs, we show that the references recommended by GPT-4 are embedded in the relevant citation context, suggesting an even deeper conceptual internalization of the citation networks. While LLMs can aid in citation generation, we find they may also amplify existing biases and introduce new ones, potentially skewing scientific discourse. Our results underscore the need for identifying the model’s biases and for developing balanced methods to interact with LLMs in general [25].
Our setting differs from previous work which either let the LLM generate short papers or literature reviews, or is prompted for the most important papers on a certain topic [35]. We argue that these methods are more susceptible to the LLM’s knowledge cut-off [11, 13, 26, 32, 33]. Moreover, the evaluation of the suggested references mostly focuses on their existence, bibliometric accuracy, or qualitative judgement by domain experts [27]. Finally, another strand of the literature focuses on improving LLMs via search and retrieval-augmented generation [21] or to reduce their hallucination rate via self-consistency [2] to enhance their capabilities in systematic literature reviews [31].
2 Generating Citations with Large Language Models
Our data consists of $166$ papers published at AAAI ($25$), NeurIPS ($72$), ICML ($38$), and ICLR ($31$) for a total of $3,066$ references. Our data collection process is depicted in Figure 1 (see Appendix A for more details) and begins by retrieving relevant papers from arXiv, focusing on those within the machine learning category (cs.LG) and posted between March 2022 and October 2023 (after GPT-4-0613’s knowledge cut-off date). We ensure these papers reference the target conferences and exclude those that do not meet our criteria using specific keywords (e.g., workshop). The papers are also verified on Semantic Scholar where we store additional metadata, such as all the reference titles with corresponding Semantic Scholar IDs to construct the citation networks (see Appendix Table LABEL:tab:paper_details for a full list of all the included papers).
We split the main content, which includes the author information, conference information, abstract, and introduction, from the ground truth references. Next, we prompt GPT-4 with the main content accompanied by the default system message: “You are a helpful assistant” and the following prompt:
Below, we share with you a written introduction to a paper and have omitted the references. Numbers between square brackets indicate citations. Can you give us a suggestion for an explicit reference associated with each number? Do not return anything except the citation number between square brackets and the corresponding reference. $=$ [main content]
We then post-process GPT-4’s response to extract the title, venue, publication year, author names, and number of authors for each generated reference (see Appendix A for more details). To assess the robustness of this approach, we repeat this “vanilla” approach five times for all $166$ papers.
A well-known issue in text generation by LLMs are hallucinations or confabulations, which refer to generated content that is nonsensical or untruthful in relation to certain sources, i.e., factual mistakes about historical events [43]. This is particularly problematic for the generation of scholarly references, as LLMs can fabricate references that do not exist or introduce subtle errors, making it impossible to retrieve the actual references [35]. There are two main approaches to verify the existence of LLM-generated references: one involves asking additional questions to the LLM to verify its self-consistency [2], and the second approach utilizes external databases to verify a reference’s existence [7]. In our experiment, we opt for the latter and determine via title and author names matching with Semantic Scholar entries whether the generated references exist (see Appendix A for more details). Finally, we also build on our “vanilla” approach, by introducing an “iterative” approach where we continue to prompt GPT-4 and indicate which generated references do not exist and ask to replace those with existing ones (see Appendix A for more details). The previously existing generated and the newly generated references are then merged.
In Table 1, we report the summary statistics for each of the five vanilla (and iterative: results between brackets) runs. On average, $65\%$ ($86\%$) of the generated references match with an entry in Semantic Scholar, while $13\%$ ($14\%$) and $17\%$ ($20\%$) of them appear in the introduction or paper itself, respectively. We further show that about $7\%$ ($7\%$) of the generated and ground truth references match pairwise. However, this number grows to $13\%$ ($14\%$) if we only consider the uniquely identifiable references (i.e., omitting references included in [4–8] as there is no one-to-one correspondence).
Table 1: Summary statistics (numbers in % with respect to total number of references).
Vanilla (Iterative) Run 1 Run 2 Run 3 Run 4 Run 5
Existence 64.3 63.3 62.8 64.2 67.6
(87.0) (85.5) (88.0) (86.8) (86.3)
Cited in paper 17.5 17.1 15.7 16.8 18.0
(20.0) (20.1) (18.4) (19.2) (20.8)
Cited in introduction 13.4 13.2 12.2 12.9 13.9
(14.5) (15.0) (13.5) (14.3) (15.3)
Pairwise match for 7.0 7.2 6.3 6.9 6.7
all references (7.1) (7.3) (6.6) (7.0) (7.1)
Pairwise match for uniquely 12.5 13.7 12.5 13.7 13.3
identifiable references (12.5) (14.1) (12.9) (13.7) (14.0)
3 Reflecting Human Citation Patterns
Figure 2 displays the characteristics of the ground truth and generated references, and separately the characteristics of the generated references which match with a Semantic Scholar entry, and those which do not exist according to this database. Overall, we observe a remarkable similarity between human and LLM citation patterns and a large consistency between GPT-4’s existing and non-existent generated references, indicating the model’s internalization of citation patterns. In Appendix Figure B1, we also show that the newly generated papers from the “iterative” approach show nearly identical distributions.
The distributions of the title lengths show that the existing generated reference titles tend to be the shortest, while the non-existent generated reference titles are more similar in length to the ground truth, which indicates a learned pattern. Overall, the first effect dominates, so the average is skewed towards shorter titles for generated references (Figure 2b). The temporal analysis reveals a similar pattern where the non-existent generated references follow a distribution that is more similar to the ground truth than the existent ones (Figure 2c). The existent generated references are on average less recent, which might be understood as follows: the sources from which the references can be learned (before the knowledge cut-off) must be more recent than the source itself, and this takes some time.
The distribution of the number of authors highlights a notable difference, with ground truth references typically involving three authors versus two for generated references, though the frequent use of “et al.” in the generated references complicates exact author counts (Figure 2d). To further examine the potential impact of the “et al.” problem, we only consider the existing generated references and their ground truth counterpart in Appendix Figure B2. There, we compare the characteristics of the references between two data sources, namely the original source (the paper or GPT generation) and the available information on Semantic Scholar. The similarity between the distributions of all characteristics shows that the data source has no impact and “et al.” does not cause this observation.
The publication venue distributions show that for most venues the ground truth has the highest relative representation, followed closely by existing generated references, with non-existent generated references displaying the largest proportion of “Others” (Figure 2e). In Appendix Figure B4, we observe that the distributions of publication venues for both ground truth and generated references are very similar across the various conferences, i.e., AAAI, NeurIPS, ICML, and ICLR. The pairwise transition matrix from ground truth to generated publication venues at the reference level indicates a large overall agreement, but with a strong preference in GPT-4 generated references for arXiv, NeurIPS, and “Others” in the case of disagreement. The preference for NeurIPS may be due to the relatively large number of NeurIPS papers in our sample and the large share of arXiv and “Others” points to favoring a wider array of venues which may potentially dilute the perceived relevance of key conferences. Finally, the scatter plot affirms the strong pairwise correlation between the ground truth and generated references to the top conferences at the individual paper level.
Most prominently, we observe a significant citation bias in the existing generated references, which have a median citation count of $1,326$ higher than ground truth references (Figure 2f). In Appendix Figure B3, we compare the characteristics for the corresponding ground truth references of existing and non-existent references, and for the existing references which also appear in the paper itself. We observe that the ground truth papers which correspond to existing references that appear in the paper itself have by far the most citations, followed by the existing references, and the ground truth papers corresponding to non-existent references have the lowest numbers of citations. These findings further indicate the tendency for GPT-4 to more easily generate references to highly cited papers. Finally, the distribution of references indicates that ground truth references cite slightly more papers than existing generated references, with a median difference of $6$ (Figure 2g).
Figure 2: Properties of the ground truth and generated introduction references for the vanilla strategy. This figure displays the properties of the ground truth ($n=14,554$, in blue) and generated references ($n=14,554$, in green), further subdividing the generated references into existing ($n=9,376$, in orange) and non-existent ($n=5,178$, in red), from the original data sources of five runs for the vanilla strategy with GPT-4. a, The average percentage of existing generated references in total ($64.4\%$) and for each publication venue under the vanilla and iterative strategy, with dots representing the percentage for each of the five runs with GPT-4. b, The distribution of the number of characters in the title shows some differences between the ground truth (median $62$) and generated (median $58$) references with the non-existent being slightly longer and with a larger variance compared to the existing generations. c, The distribution over time reveals that ground truth references are relatively more recent than generated references, with most references post-$2010$. The temporal distribution of the non-existent generated references aligns more with the ground truth than the existing generated references. d, The distribution of the number of authors demonstrates a disparity between the ground truth and generated references, having median values of three and two, respectively. However, GPT-4 more often generates “et al.” which does not allow for an exact computation, especially for the non-existent references. e, The distribution of publication venues shows that for most venues the ground truth has the highest relative representation, followed closely by existing references. The non-existent references deviate more from the ground truth as the proportion of “Others” is substantially larger. f, The distributions of citations for ground truth and existing generated references reveal a substantial citation bias in the generated references with a difference in median citations of $1,326$. g, Finally, the distribution of references shows that ground truth references cite slightly more papers than the existing generated references with a median difference in median references of $6$.
Figure 3: The citation bias in existing generated references is not due to the recency of ground truth references. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher citation count compared to their corresponding ground truth ($n=9,376$, in blue) across subperiods. a, The citation counts across time for the ground truth and existing generated references reveal that the most recent references have a relatively low number of citations. The difference in median citations between the existing generated references and their corresponding ground truth references is 1,257. Since the ground truth references are relatively more recent compared to the existing generated references, we examine whether the observed citation bias is related to the recency of ground truth references. b, The distributions of citations by subperiod reveal that the existing generated references consistently exhibit a higher citation count than their corresponding ground truth counterparts. c, The difference in median citations is most pronounced in the early and late subperiods, i.e., $\leq 1988$, 2010-2016, and 2017-2023.
Figure 4: The citation bias in existing generated references is consistent across title length, number of authors, and publication venue. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher citation count compared to their corresponding ground truth ($n=9,376$, in blue) across title length, number of authors, and publication venue. a, The citation counts across the number of characters in title reveals that the discrepancy in number of citations between the existing generated and ground truth references is consistent over various title lengths. b and c, The distributions of citation counts per number of authors and publication venues show that the existing generated references consistently exhibit a higher citation count than their corresponding ground truth counterparts.
Figure 5: The generated references display similar citation network properties as the ground truth references but with a heightened citation bias. This figure displays how the existing generated references ($n=2945$, first run of vanilla strategy) are embedded in the citation network of the focal papers ($m=166$ in total). a, We depict the connections between the focal paper, the ground truth references, and the existing generated references by showing the underlying citation graphs. An arrow from A to B indicates that A cites B. We identify the focal paper (in blue), generated references that appear in the introduction (in green) or in the paper (in yellow), generated references that are linked to ground truth or other generated references (in orange), generated references that are completely isolated (in purple), and ground truth references that are not cited by GPT-4 (in gray). b, The majority of generated references does not appear in the introduction or paper itself, but is somehow connected to the ground truth references as only a small fraction of generated references is completely isolated. c, The heightened citation bias is most emphasized for generated references that appear in the introduction or the paper, with isolated generated references having the lowest number of citations. d, The normalized average clustering coefficients of the ground truth (green and grey nodes) and the existing generated references (green, yellow, orange, and purple nodes) indicate that GPT-4’s internalization of citation patterns extends to citation network properties. The clustering coefficient for a node A is given by #triangles through A#possible triangles through $\frac{\#\text{triangles through A}}{\#\text{possible triangles through A}}$. The average is computed across the coefficients of all nodes in the respective graph (excl. nodes with coefficient zero) and indicates the tendency of the respective references to appear in clusters. e, The non-isolated generated references are tightly connected to the ground truth references, both on an individual level (Boolean edge density) as well as an aggregate level (edge expansion). The Boolean edge density is the fraction of non-isolated generations (orange nodes) that are connected to at least one ground truth reference (green and grey nodes) per focal paper. The edge expansion between those two sets is defined as the number of edges between the two sets divided by the smallest set size. f, The number of references is similar across all categories, except for the isolated generated references which have substantially less references.
4 Heightened Citation Bias
Figure 3 demonstrates that the citation bias observed in GPT-4 generated references is not merely a consequence of the recency of ground truth references. Specifically, the existing generated references show consistently higher citation counts compared to their ground truth counterparts across various subperiods. Figure 3a illustrates that ground truth references, particularly the most recent ones, tend to have lower citation counts. Despite the ground truth references being more recent on average, the citation counts of existing generated references remain significantly higher. Figure 3b further breaks down the citation distributions by subperiods, reaffirming that generated references consistently have higher citation counts than their corresponding ground truth references. Figure 3c highlights that this citation discrepancy is most pronounced in both the earliest ($\leq$1988) and the most recent (2010-2016 and 2017-2023) subperiods, indicating that the citation bias persists across different time frames.
In Figure 4, we find that the heightened citation bias in generated references remains also after controlling for other possible confounding factors, such as title length, number of authors, and publication venue. Figure 4a shows that the citation discrepancy is consistent across different title lengths, with generated references tending to be more highly cited regardless of title length variations. Figure 4(b,c) examine the distributions of citation counts by the number of authors and publication venues, respectively. The data indicates that, regardless of the number of authors or the publication venue, generated references have higher citation counts compared to the ground truth references. In Appendix Figure B5 and Appendix Figure B6, we confirm that our findings are robust for the influential citation count which can be retrieved from Semantic Scholar [34]. This consistency across multiple factors underscores the inherent bias of LLMs towards generating references to highly cited papers, irrespective of other characteristics of the references.
5 Large Language Models and Human Citation Networks
Figure 5 displays the properties of the ground truth and generated citation networks. In Figure 5a, we identify the focal paper (in blue), generated references that appear in the introduction (in green) or later in the paper (in yellow), generated references that are linked to ground truth or other generated references (in orange), generated references that are completely isolated (in purple), and ground truth references that are not cited by GPT-4 (in gray). The majority of generated references ($+50\%$) is non-isolated, i.e., linked to the ground truth or generated references but not present in the focal paper itself, followed by a substantial amount of generated references appearing in the introduction and only a small fraction that do not appear in the introduction but still within the focal paper (Figure 5b). The remainder of generated references is completely isolated from the citation network. If GPT-4 did not pick up on human citation patterns, the generated citation network would resemble a random network containing only isolated citations. The heightened citation bias is also most pronounced for references that appear within the introduction or paper, with isolated generated references having the lowest number of citations (Figure 5c). This finding further indicates the tendency for GPT-4 to more easily identify and generate references to highly cited papers. The number of references is similar across all categories, except for the isolated generated references which have substantially less references (Figure 5f).
[…]
6 Discussion
We designed an experiment to explore the intrinsic citation behavior of GPT-4 and its potential biases when generating scholarly references. While previous work focuses on LLMs generating short papers or literature reviews [27, 35], we let GPT-4 generate suggestions of scholarly references for anonymized in-text citations. Importantly, we do not enhance the LLM through search and retrieval-augmented generation, but evaluate the model’s internalization of citation patterns in its parametric knowledge encapsulated during its (pre-)training.
Our findings are significant as they represent a first step towards understanding the real-world impact of LLMs in scientific research [3, 44]. By highlighting the heightened citation bias in GPT-4 generated references, we demonstrate the model’s tendency to favor highly cited papers, which could exacerbate existing biases in scientific discourse. This evaluation moves beyond more traditional LLM benchmarks [30], emphasizing the practical implications of deploying these models in academic contexts [12]. The results suggest that while LLMs have the potential to streamline various aspects of research, careful consideration is needed to mitigate the amplification of biases.
One plausible hypothesis for the heightened citation bias observed in GPT-4 is the increased frequency of citations to heavily cited papers within the model’s training data. This prevalence makes these references more likely to be generated accurately and recognized as existent. Additionally, such biases may stem from generic training effects, where models preferentially learn patterns that are more common in the data, leading to biases towards shorter titles, more heavily cited, and slightly less recent works. These tendencies may persist despite improvements in data quantity or model sophistication [15, 19].
We developed and open-source an extensible, automated pipeline to systematically analyze the references generated by LLMs. Although our methodology is robust, it is not without limitations. The use of simple prompts and the zero-shot setting [17] aims to minimize bias in the generation process, but this simplicity might not capture the full spectrum of potential LLM capabilities. There are numerous alternative approaches and prompt designs that future research can explore to enhance the accuracy and relevance of generated references [37, 39, 42]. However, our iterative approach already indicates that biases remain inherent in these generations.
In conclusion, while LLMs like GPT-4 can significantly aid in citation generation, they also risk amplifying existing biases and introducing new ones, potentially skewing the structuring and the dissemination of scientific knowledge. Our study underscores the necessity for developing balanced methods to interact with LLMs, incorporating diverse datasets, and implementing bias mitigation strategies. Fair prompting techniques [22], for instance, can be employed to reduce bias, but continuous vigilance and methodological innovation are required to ensure that the integration of LLMs into academic workflows promotes accurate knowledge dissemination.
Acknowledgments and Disclosure of Funding
Andres Algaba acknowledges a fellowship from the Research Foundation Flanders under Grant No.1286924N. Vincent Ginis acknowledges support from Research Foundation Flanders under Grant No.G032822N and G0K9322N.
Appendix A Data
We describe the steps of our automated pipeline to retrieve all the necessary information for our analysis. Our data collection resulted in $166$ papers published at AAAI ($25$), NeurIPS ($72$), ICML ($38$), and ICLR ($31$) for a total of 3,066 references (see Appendix Table LABEL:tab:paper_details for a full list of included papers). The data collection pipeline uses GPT-4 to postprocess parts of the data, which costs approximately 14 dollars for our experiment. Note that these steps only have to be carried out once for the data collection. However, steps 4 and 5 are also used to postprocess and enrich the information of the generated references and will need to be carried out for each run. The experiment was run on 4 November 2023 and each step was manually verified and tested.
Step 1. ArXiv
We search for all papers on arXiv originally posted between 1 March 2022 and 31 October 2023 in the machine learning (cs.LG) category which refer to AAAI, NeurIPS, ICLR, or ICML in their journal reference. Note that we also verify whether we can use all these arXiv papers given their data licenses and attribute their participation in Appendix Table LABEL:tab:paper_details. We use keywords (i.e., workshop, tiny paper, 2020, 2021, track on datasets and benchmarks, and bridge) to remove papers that do not appear in the conference proceedings or earlier than 2022. We download and unzip the tar.gz file provided by the authors to arXiv and check whether the paper exists on Semantic Scholar via title matching. We store the title, ID, and date from arXiv and Semantic Scholar. Additionally, we store all the reference titles with their corresponding ID from Semantic Scholar [16].
Step 2. Tex
[…]
Step 3. PDF
We transform the PDF to txt and split the main content of the paper (author information, conference information, abstract, and introduction) from the references. We then look for all in-text citations by using a regex pattern to capture numbers in between square brackets and match them with the reference list. This approach ensures that we only keep references that are cited in the introduction. We store the main content of the paper and the references cited in the introduction in separate txt files.
Step 4. Postprocessing
A large number of variations and inconsistencies in the reference lists makes it difficult to structurally extract and analyze all the author information, title, publication venue, and year. We noticed that this behavior was even more outspoken in the LLM-generated references. Therefore, we examine the capabilities of GPT-4 to impose a structure on the reference list by postprocessing the data. We feed GPT-4 the reference list in txt accompanied by the default system message: “You are a helpful assistant” and the following postprocessing prompt:
Below, we share with you a list of references with their corresponding citation number between square brackets. Could you for each reference extract the authors, the number of authors, title, publication year, and publication venue? Please only return the extracted information in a markdown table with the citation number (without brackets), authors, number of authors, title, publication year, and publication venue as columns. $=$ [GPT-4 generated reference list]
We then store the markdown table in a csv. GPT-4 successfully structures the information and makes it more consistent, for example, by removing syllable hyphens. Sometimes a small hick-up is introduced (e.g., adding a final row with “…”), but these are manually solved in the verification process. Note that we also prompt for the number of authors. While we can easily compute the number of authors via the meta-data from Semantic Scholar, it allows us to verify the accuracy of GPT-4 on this task as we will use it later on to postprocess the generated references where a ground truth may be unavailable.
Step 5. Semantic Scholar
We enrich the information from the introduction references by matching the extracted title from the csv file in the previous step with the reference titles that we extracted from Semantic Scholar in step 1. This approach provides an extra check that GPT-4 does not change the title information in Step 4. After matching, we can use the Semantic Scholar ID to retrieve the publication venue, year, authors, citation count, influential citation count, and reference count [16]. Additionally, we store the IDs of the papers to which the introduction references themselves refer.
Step 6. “Vanilla” prompting
We prompt GPT-4 with the main content, which includes the author information, conference information, abstract, and introduction, accompanied by the default system message: “You are a helpful assistant” and the following prompt:
Below, we share with you a written introduction to a paper and have omitted the references. Numbers between square brackets indicate citations. Can you give us a suggestion for an explicit reference associated with each number? Do not return anything except the citation number between square brackets and the corresponding reference. $=$ [main content]
We then post-process GPT-4’s response to extract the title, venue, publication year, author names, and number of authors for each generated reference using the same approach as described in step 4. We repeat this “vanilla” approach five times for all $166$ papers.
Step 7. Existence check
We determine whether the generated references exist via title and author names matching with Semantic Scholar entries [16]. We search Semantic Scholar for the three best matches based on the reference’s title and then compute the title and author names similarity. For titles, we measure the similarity between the Semantic Scholar match and the generated reference by comparing the best matching substring. For authors, we compare them by splitting into tokens (words), removing duplicates, and then calculating the similarity based on the best partial match of the sets of tokens. In case of “et al.,” we only consider the first author. The similarity is computed by character-level comparison. We determined the thresholds for the title and authors scores by manually labelling $100$ matches as true or false and minimizing the false positive rate. We obtain on this sample an accuracy of $95\%$ with $5$ false positives, i.e. generated references falsely classified as non-existent.
Step 8. “Iterative” prompting
We also build on our “vanilla” approach, by introducing an “iterative” approach where we prompt GPT-4 with the main content accompanied by the default system message: “You are a helpful assistant” and the following prompt:
[vanilla prompt + GPT-4’s response ] The following references associated with these citation numbers: [numbers of non-existent generated references ] do not exist. Can you replace all these non-existent references with existing ones? Keep the other references as they are. Do not return anything except the citation number between square brackets and the corresponding reference. $=$ [main content]
We again postprocess GPT-4’s response using the same approach as described in steps 4, 5, and 7. The previously existing generated and the newly generated references are then merged.
Appendix B Additional Figures
Figure B1: Properties of the ground truth and generated introduction references for the iterative strategy are consistent with the properties of the vanilla strategy. This figure displays the properties of the ground truth ($n=5,178$, in blue) and generated references ($n=5,178$, in green), further subdividing the generated references into existing ($n=3,244$, in orange) and non-existent categories ($n=1,934$, in red), from the original data sources of five runs for the iterative strategy with GPT-4. Note that these are the references which are labelled “non-existent” in the vanilla strategy. a, b, c, d, e, f and g, The iterative results exhibit very similar properties to the vanilla results shown in Fig. 1.
Figure B2: The properties of the existing generated references and their corresponding ground truth are consistent between the original data sources and Semantic Scholar data. This figure compares the computation of the properties of the existing generated references ($n=9.376$) and their corresponding ground truth ($n=9.376$) between the data from the original sources (in dark blue and orange, as shown in Fig. 1) and from Semantic Scholar (in light blue and orange). a, The distributions of the number of characters in title for the existing generated references and their corresponding ground truth are very similar between the data from the original sources and Semantic Scholar. b, There is a discrepancy between the data from the original sources and Semantic Scholar for the number of authors in the existing generated references due to the extensive use of “et al”. This discrepancy results in a relatively larger portion of three authors or more, but does not change the previous conclusions. c, The distributions over time are very similar between the data from the original sources and Semantic Scholar. d, There is a discrepancy between the data from the original sources and Semantic Scholar for the publication venues. The discrepancy is consistent across the existing generated references and their corresponding ground truth as both have a lower number of arXiv papers and a larger number of ICLR papers.
Figure B3: Ground truth papers which correspond to existing generated references that appear in the paper have substantially more citations. This figure displays the properties of the ground truth references which correspond to existing generated references ($n=9,376$, in yellow), the subset of existing generated references which appear in the paper itself ($n=2,474$, in blue), and the non-existent generated references ($n=5,178$, in green), from the original data sources of five runs for the vanilla strategy with GPT-4. a, b, c, d, e and f, The ground truth papers which correspond to existing references which appear in the paper have by far the most citations, followed by the existing references, and the ground truth papers corresponding to non-existent references have the lowest numbers of citations. These findings further indicate the tendency for LLMs to more easily generate references to highly cited papers. The distributions of all other characteristics are very similar.
Figure B4: A high consistency in publication venue distributions between ground truth and GPT-4 generated references with a notable bias towards arXiv, NeurIPS, and “Others”. This figure displays the distributions and pairwise transition of publication venues for ground truth and generated references at the conference and individual paper and reference level. a and b, We observe that the distributions of publication venues for both ground truth and generated references are very similar across the various conferences, i.e., AAAI, NeurIPS, ICML, and ICLR. c and d, The pairwise transition matrix from ground truth to generated publication venues at the reference level indicates a large overall agreement, but with a strong preference in GPT-4 generated references for arXiv, NeurIPS, and “Others” in the case of disagreement. The preference for NeurIPS may be due to the relatively large number of NeurIPS papers in our sample and the large share of arXiv and “Others” points to favoring a wider array of venues which may potentially dilute the perceived relevance of key conferences. e, The scatter plot shows for each paper the number of ground truth references to one of the top conferences, i.e., AAAI, NeurIPS, ICML, and ICLR, and the corresponding number of generated references (×$\times 5$ for five runs) which refer to the same conference.The strong pairwise correlation between the ground truth and generated references to the top conferences at the individual paper level affirms the high consistency in publication venue distributions between ground truth and GPT-4 generated references.
Figure B5: The influential citation bias in existing generated references is unrelated to the recency of ground truth references. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher influential citation count compared to their corresponding ground truth ($n=9,376$, in blue) across subperiods. a, b and c, Note that the influential citation count is retrieved from Semantic Scholar [34].
Figure B6: The infuential citation bias in existing generated references is unrelated to title length, number of authors, and publication venue. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher influential citation count compared to their corresponding ground truth ($n=9,376$, in blue) across title length, number of authors, and publication venue. a, b and c, Note that the influential citation count is retrieved from Semantic Scholar [34].
Appendix C Additional Tables
Table C1: Papers included in the analysis.
Conference Authors Title
AAAI Jakob Weissteiner, Jakob Heiss, Julien Siems, Sven Seuken Bayesian Optimization-based Combinatorial Assignment
AAAI Gobinda Saha, Kaushik Roy Continual Learning with Scaled Gradient Projection
AAAI Ruizhe Zheng, Jun Li, Yi Wang, Tian Luo, Yuguo Yu ScatterFormer: Locally-Invariant Scattering Transformer for Patient-Independent Multispectral Detection of Epileptiform Discharges
AAAI Sahil Manchanda, Sayan Ranu Lifelong Learning for Neural powered Mixed Integer Programming
AAAI Joris Guérin, Kevin Delmas, Raul Sena Ferreira, Jérémie Guiochet Out-Of-Distribution Detection Is Not All You Need
AAAI Taha Belkhouja, Yan Yan, Janardhan Rao Doppa Training Robust Deep Models for Time-Series Domain: Novel Algorithms and Theoretical Analysis
AAAI Minsoo Kang, Suhyun Kim GuidedMixup: An Efficient Mixup Strategy Guided by Saliency Maps
AAAI Su Kim, Dongha Lee, SeongKu Kang, Seonghyeon Lee, Hwanjo Yu Learning Topology-Specific Experts for Molecular Property Prediction
AAAI Daniel Silver, Tirthak Patel, Devesh Tiwari QUILT: Effective Multi-Class Classification on Quantum Computers Using an Ensemble of Diverse Quantum Classifiers
AAAI Kevin Osanlou, Jeremy Frank, Andrei Bursuc, Tristan Cazenave, Eric Jacopin, Christophe Guettier, J. Benton Solving Disjunctive Temporal Networks with Uncertainty under Restricted Time-Based Controllability using Tree Search and Graph Neural Networks
AAAI Joar Skalse, Alessandro Abate Misspecification in Inverse Reinforcement Learning
AAAI Edward Ayers, Jonathan Sadeghi, John Redford, Romain Mueller, Puneet K. Dokania Query-based Hard-Image Retrieval for Object Detection at Test Time
AAAI Shubham Gupta, Sahil Manchanda, Srikanta Bedathur, Sayan Ranu TIGGER: Scalable Generative Modelling for Temporal Interaction Graphs
AAAI Fanchen Bu, Dong Eui Chang Feedback Gradient Descent: Efficient and Stable Optimization with Orthogonality for DNNs
AAAI Shota Saito Hypergraph Modeling via Spectral Embedding Connection: Hypergraph Cut, Weighted Kernel $k$-means, and Heat Kernel
AAAI Haoran Luo, Haihong E, Ling Tan, Gengxian Zhou, Tianyu Yao, Kaiyang Wan DHGE: Dual-View Hyper-Relational Knowledge Graph Embedding for Link Prediction and Entity Typing
AAAI Yujin Kim, Dogyun Park, Dohee Kim, Suhyun Kim NaturalInversion: Data-Free Image Synthesis Improving Real-World Consistency
AAAI Tairan He, Weiye Zhao, Changliu Liu AutoCost: Evolving Intrinsic Cost for Zero-violation Reinforcement Learning
AAAI Shijie Liu, Andrew C. Cullen, Paul Montague, Sarah M. Erfani, Benjamin I. P. Rubinstein Enhancing the Antidote: Improved Pointwise Certifications against Poisoning Attacks
AAAI Fan Zhou, Chen Pan, Lintao Ma, Yu Liu, Shiyu Wang, James Zhang, Xinxin Zhu, Xuanwei Hu, Yunhua Hu, Yangfei Zheng, Lei Lei, Yun Hu SLOTH: Structured Learning and Task-based Optimization for Time Series Forecasting on Hierarchies
AAAI Christopher W. F. Parsonson, Alexandre Laterre, Thomas D. Barrett Reinforcement Learning for Branch-and-Bound Optimisation using Retrospective Trajectories
AAAI Sourya Basu, Prasanna Sattigeri, Karthikeyan Natesan Ramamurthy, Vijil Chenthamarakshan, Kush R. Varshney, Lav R. Varshney, Payel Das Equi-Tuning: Group Equivariant Fine-Tuning of Pretrained Models
AAAI Harry Rubin-Falcone, Joyce Lee, Jenna Wiens Forecasting with Sparse but Informative Variables: A Case Study in Predicting Blood Glucose
AAAI Pierre Le Pelletier de Woillemont, Rémi Labory, Vincent Corruble Automated Play-Testing Through RL Based Human-Like Play-Styles Generation
AAAI Kai Klede, Leo Schwinn, Dario Zanca, Björn Eskofier FastAMI – a Monte Carlo Approach to the Adjustment for Chance in Clustering Comparison Metrics
NeurIPS Dhananjay Bhaskar, Kincaid MacDonald, Oluwadamilola Fasina, Dawson Thomas, Bastian Rieck, Ian Adelstein, Smita Krishnaswamy Diffusion Curvature for Estimating Local Curvature in High Dimensional Data
NeurIPS Shiro Takagi On the Effect of Pre-training for Transformer in Different Modality on Offline Reinforcement Learning
NeurIPS Yue Yu, Yuchen Zhuang, Jieyu Zhang, Yu Meng, Alexander Ratner, Ranjay Krishna, Jiaming Shen, Chao Zhang Large Language Model as Attributed Training Data Generator: A Tale of Diversity and Bias
NeurIPS Lingfeng Sun, Haichao Zhang, Wei Xu, Masayoshi Tomizuka PaCo: Parameter-Compositional Multi-Task Reinforcement Learning
NeurIPS Yang Yue, Rui Lu, Bingyi Kang, Shiji Song, Gao Huang Understanding, Predicting and Better Resolving Q-Value Divergence in Offline-RL
NeurIPS Jiaqi Leng, Yuxiang Peng, Yi-Ling Qiao, Ming Lin, Xiaodi Wu Differentiable Analog Quantum Computing for Optimization and Control
NeurIPS Kyriakos Flouris, Ender Konukoglu Canonical normalizing flows for manifold learning
NeurIPS Yuchen Bai, Jean-Baptiste Durand, Florence Forbes, Grégoire Vincent Semantic segmentation of sparse irregular point clouds for leaf wood discrimination
NeurIPS Lorenzo Giambagli, Lorenzo Buffoni, Lorenzo Chicchi, Duccio Fanelli How a student becomes a teacher: learning and forgetting through Spectral methods
NeurIPS Hanbyul Lee, Qifan Song, Jean Honorio Support Recovery in Sparse PCA with Incomplete Data
NeurIPS Zhang-Wei Hong, Aviral Kumar, Sathwik Karnik, Abhishek Bhandwaldar, Akash Srivastava, Joni Pajarinen, Romain Laroche, Abhishek Gupta, Pulkit Agrawal Beyond Uniform Sampling: Offline Reinforcement Learning with Imbalanced Datasets
NeurIPS Xiang Zhang, Ziyuan Zhao, Theodoros Tsiligkaridis, Marinka Zitnik Self-Supervised Contrastive Pre-Training For Time Series via Time-Frequency Consistency
NeurIPS Antonin Schrab, Ilmun Kim, Benjamin Guedj, Arthur Gretton Efficient Aggregated Kernel Tests using Incomplete $U$-statistics
NeurIPS Wanyun Cui, Xingran Chen Instance-based Learning for Knowledge Base Completion
NeurIPS Aurelien Lucchi, Frank Proske, Antonio Orvieto, Francis Bach, Hans Kersting On the Theoretical Properties of Noise Correlation in Stochastic Optimization
NeurIPS Minsik Cho, Saurabh Adya, Devang Naik PDP: Parameter-free Differentiable Pruning is All You Need
NeurIPS Guangxi Li, Ruilin Ye, Xuanqiang Zhao, Xin Wang Concentration of Data Encoding in Parameterized Quantum Circuits
NeurIPS Xinrui Wang, Wenhai Wan, Chuanxin Geng, Shaoyuan LI, Songcan Chen Beyond Myopia: Learning from Positive and Unlabeled Data through Holistic Predictive Trends
NeurIPS Zihan Liu, Yun Luo, Lirong Wu, Zicheng Liu, Stan Z. Li Towards Reasonable Budget Allocation in Untargeted Graph Structure Attacks via Gradient Debias
NeurIPS Dingfan Chen, Raouf Kerkouche, Mario Fritz Private Set Generation with Discriminative Information
NeurIPS Zhan Yu, Hongshun Yao, Mujin Li, Xin Wang Power and limitations of single-qubit native quantum neural networks
NeurIPS Ibrahim Alabdulmohsin, Xiaohua Zhai, Alexander Kolesnikov, Lucas Beyer Getting ViT in Shape: Scaling Laws for Compute-Optimal Model Design
NeurIPS Manzil Zaheer, Kenneth Marino, Will Grathwohl, John Schultz, Wendy Shang, Sheila Babayan, Arun Ahuja, Ishita Dasgupta, Christine Kaeser-Chen, Rob Fergus Learning to Navigate Wikipedia by Taking Random Walks
NeurIPS Dohyun Kwon, Ying Fan, Kangwook Lee Score-based Generative Modeling Secretly Minimizes the Wasserstein Distance
NeurIPS Zhaoqi Li, Lillian Ratliff, Houssam Nassif, Kevin Jamieson, Lalit Jain Instance-optimal PAC Algorithms for Contextual Bandits
NeurIPS Masaki Adachi, Satoshi Hayakawa, Martin Jørgensen, Harald Oberhauser, Michael A. Osborne Fast Bayesian Inference with Batch Bayesian Quadrature via Kernel Recombination
NeurIPS Zhiqin Yang, Yonggang Zhang, Yu Zheng, Xinmei Tian, Hao Peng, Tongliang Liu, Bo Han FedFed: Feature Distillation against Data Heterogeneity in Federated Learning
NeurIPS Daniel Vial, Sujay Sanghavi, Sanjay Shakkottai, R. Srikant Minimax Regret for Cascading Bandits
NeurIPS Fabian Zaiser, Andrzej S. Murawski, Luke Ong Exact Bayesian Inference on Discrete Models via Probability Generating Functions: A Probabilistic Programming Approach
NeurIPS Cheng Chi, Amine Mohamed Aboussalah, Elias B. Khalil, Juyoung Wang, Zoha Sherkat-Masoumi A Deep Reinforcement Learning Framework For Column Generation
NeurIPS Mathieu Molina, Patrick Loiseau Bounding and Approximating Intersectional Fairness through Marginal Fairness
NeurIPS Shuai Zhang, Hongkang Li, Meng Wang, Miao Liu, Pin-Yu Chen, Songtao Lu, Sijia Liu, Keerthiram Murugesan, Subhajit Chaudhury On the Convergence and Sample Complexity Analysis of Deep Q-Networks with $\varepsilon$-Greedy Exploration
NeurIPS Changlong Wu, Mohsen Heidari, Ananth Grama, Wojciech Szpankowski Precise Regret Bounds for Log-loss via a Truncated Bayesian Algorithm
NeurIPS Ching-Yao Chuang, Stefanie Jegelka Tree Mover’s Distance: Bridging Graph Metrics and Stability of Graph Neural Networks
NeurIPS Felix Biggs, Antonin Schrab, Arthur Gretton MMD-FUSE: Learning and Combining Kernels for Two-Sample Testing Without Data Splitting
NeurIPS Thomas Fel, Victor Boutin, Mazda Moayeri, Rémi Cadène, Louis Bethune, Léo andéol, Mathieu Chalvidal, Thomas Serre A Holistic Approach to Unifying Automatic Concept Extraction and Concept Importance Estimation
NeurIPS Manel Baradad, Chun-Fu Chen, Jonas Wulff, Tongzhou Wang, Rogerio Feris, Antonio Torralba, Phillip Isola Procedural Image Programs for Representation Learning
NeurIPS Yang Ni Bivariate Causal Discovery for Categorical Data via Classification with Optimal Label Permutation
NeurIPS Gauthier Guinet, Saurabh Amin, Patrick Jaillet Effective Dimension in Bandit Problems under Censorship
NeurIPS Kyungmin Lee, Jinwoo Shin RenyiCL: Contrastive Representation Learning with Skew Renyi Divergence
NeurIPS Yihe Wang, Yu Han, Haishuai Wang, Xiang Zhang Contrast Everything: A Hierarchical Contrastive Framework for Medical Time-Series
NeurIPS Artyom Sorokin, Nazar Buzun, Leonid Pugachev, Mikhail Burtsev Explain My Surprise: Learning Efficient Long-Term Memory by Predicting Uncertain Outcomes
NeurIPS Yipeng Kang, Tonghan Wang, Xiaoran Wu, Qianlan Yang, Chongjie Zhang Non-Linear Coordination Graphs
NeurIPS Niv Giladi, Shahar Gottlieb, Moran Shkolnik, Asaf Karnieli, Ron Banner, Elad Hoffer, Kfir Yehuda Levy, Daniel Soudry DropCompute: simple and more robust distributed synchronous training via compute variance reduction
NeurIPS Jack Richter-Powell, Yaron Lipman, Ricky T. Q. Chen Neural Conservation Laws: A Divergence-Free Perspective
NeurIPS Peide Huang, Mengdi Xu, Jiacheng Zhu, Laixi Shi, Fei Fang, Ding Zhao Curriculum Reinforcement Learning using Optimal Transport via Gradual Domain Adaptation
NeurIPS Mark D. McDonnell, Dong Gong, Amin Parveneh, Ehsan Abbasnejad, Anton van den Hengel RanPAC: Random Projections and Pre-trained Models for Continual Learning
NeurIPS Haoyuan Sun, Kwangjun Ahn, Christos Thrampoulidis, Navid Azizan Mirror Descent Maximizes Generalized Margin and Can Be Implemented Efficiently
NeurIPS Rui M. Castro, Fredrik Hellström, Tim van Erven Adaptive Selective Sampling for Online Prediction with Experts
NeurIPS Tonghan Wang, Paul Dütting, Dmitry Ivanov, Inbal Talgam-Cohen, David C. Parkes Deep Contract Design via Discontinuous Networks
NeurIPS Sourya Basu, Pulkit Katdare, Prasanna Sattigeri, Vijil Chenthamarakshan, Katherine Driggs-Campbell, Payel Das, Lav R. Varshney Efficient Equivariant Transfer Learning from Pretrained Models
NeurIPS Qianyi Li, Haim Sompolinsky Globally Gated Deep Linear Networks
NeurIPS Jonatha Anselmi, Bruno Gaujal, Louis-Sébastien Rebuffi Reinforcement Learning in a Birth and Death Process: Breaking the Dependence on the State Space
NeurIPS Zhanpeng Zhou, Yongyi Yang, Xiaojiang Yang, Junchi Yan, Wei Hu Going Beyond Linear Mode Connectivity: The Layerwise Linear Feature Connectivity
NeurIPS Jinyu Cai, Jicong Fan Perturbation Learning Based Anomaly Detection
NeurIPS Dan Zhao Combining Explicit and Implicit Regularization for Efficient Learning in Deep Networks
NeurIPS Leonard Papenmeier, Luigi Nardi, Matthias Poloczek Increasing the Scope as You Learn: Adaptive Bayesian Optimization in Nested Subspaces
NeurIPS Yang Song, Qiyu Kang, Sijie Wang, Zhao Kai, Wee Peng Tay On the Robustness of Graph Neural Diffusion to Topology Perturbations
NeurIPS Ibrahim Alabdulmohsin, Behnam Neyshabur, Xiaohua Zhai Revisiting Neural Scaling Laws in Language and Vision
NeurIPS Salva Rühling Cachay, Bo Zhao, Hailey Joren, Rose Yu DYffusion: A Dynamics-informed Diffusion Model for Spatiotemporal Forecasting
NeurIPS Indradyumna Roy, Soumen Chakrabarti, Abir De Maximum Common Subgraph Guided Graph Retrieval: Late and Early Interaction Networks
NeurIPS Divin Yan, Gengchen Wei, Chen Yang, Shengzhong Zhang, Zengfeng Huang Rethinking Semi-Supervised Imbalanced Node Classification from Bias-Variance Decomposition
NeurIPS Zhiying Lu, Hongtao Xie, Chuanbin Liu, Yongdong Zhang Bridging the Gap Between Vision Transformers and Convolutional Neural Networks on Small Datasets
NeurIPS Rémi Leluc, François Portier, Johan Segers, Aigerim Zhuman A Quadrature Rule combining Control Variates and Adaptive Importance Sampling
NeurIPS Kwangjun Ahn, Xiang Cheng, Hadi Daneshmand, Suvrit Sra Transformers learn to implement preconditioned gradient descent for in-context learning
NeurIPS Annie S. Chen, Archit Sharma, Sergey Levine, Chelsea Finn You Only Live Once: Single-Life Reinforcement Learning
NeurIPS Sen Lin, Daouda Sow, Kaiyi Ji, Yingbin Liang, Ness Shroff Non-Convex Bilevel Optimization with Time-Varying Objective Functions
NeurIPS Carl Hvarfner, Erik Hellsten, Frank Hutter, Luigi Nardi Self-Correcting Bayesian Optimization through Bayesian Active Learning
NeurIPS Abir De, Soumen Chakrabarti Neural Estimation of Submodular Functions with Applications to Differentiable Subset Selection
NeurIPS Ximing Lu, Sean Welleck, Jack Hessel, Liwei Jiang, Lianhui Qin, Peter West, Prithviraj Ammanabrolu, Yejin Choi Quark: Controllable Text Generation with Reinforced Unlearning
NeurIPS Weirui Ye, Pieter Abbeel, Yang Gao Spending Thinking Time Wisely: Accelerating MCTS with Virtual Expansions
NeurIPS Axel Levy, Gordon Wetzstein, Julien Martel, Frederic Poitevin, Ellen D. Zhong Amortized Inference for Heterogeneous Reconstruction in Cryo-EM
ICLR Tim Pearce, Tabish Rashid, Anssi Kanervisto, Dave Bignell, Mingfei Sun, Raluca Georgescu, Sergio Valcarcel Macua, Shan Zheng Tan, Ida Momennejad, Katja Hofmann, Sam Devlin Imitating Human Behaviour with Diffusion Models
ICLR Yi Ren, Shangmin Guo, Wonho Bae, Danica J. Sutherland How to prepare your task head for finetuning
ICLR Kieran A. Murphy, Dani S. Bassett Interpretability with full complexity by constraining feature information
ICLR Julius Adebayo, Michael Muelly, Hal Abelson, Been Kim Post hoc Explanations may be Ineffective for Detecting Unknown Spurious Correlation
ICLR Roman Levin, Valeriia Cherepanova, Avi Schwarzschild, Arpit Bansal, C. Bayan Bruss, Tom Goldstein, Andrew Gordon Wilson, Micah Goldblum Transfer Learning with Deep Tabular Models
ICLR Aviv A. Rosenberg, Sanketh Vedula, Yaniv Romano, Alex M. Bronstein Fast Nonlinear Vector Quantile Regression
ICLR Edward De Brouwer, Rahul G. Krishnan Anamnesic Neural Differential Equations with Orthogonal Polynomial Projections
ICLR Ilya Trofimov, Daniil Cherniavskii, Eduard Tulchinskii, Nikita Balabin, Evgeny Burnaev, Serguei Barannikov Learning Topology-Preserving Data Representations
ICLR Trenton Bricken, Xander Davies, Deepak Singh, Dmitry Krotov, Gabriel Kreiman Sparse Distributed Memory is a Continual Learner
ICLR Steeven Janny, Aurélien Béneteau, Madiha Nadri, Julie Digne, Nicolas Thome, Christian Wolf Eagle: Large-Scale Learning of Turbulent Fluid Dynamics with Mesh Transformers
ICLR Clement Vignac, Igor Krawczuk, Antoine Siraudin, Bohan Wang, Volkan Cevher, Pascal Frossard DiGress: Discrete Denoising diffusion for graph generation
ICLR Xinting Hu, Yulei Niu, Chunyan Miao, Xian-Sheng Hua, Hanwang Zhang On Non-Random Missing Labels in Semi-Supervised Learning
ICLR Jiefeng Chen, Timothy Nguyen, Dilan Gorur, Arslan Chaudhry Is forgetting less a good inductive bias for forward transfer?
ICLR Matthew J. Tilley, Michelle Miller, David J. Freedman Artificial Neuronal Ensembles with Learned Context Dependent Gating
ICLR Zhang-Wei Hong, Tao Chen, Yen-Chen Lin, Joni Pajarinen, Pulkit Agrawal Topological Experience Replay
ICLR Lingkai Kong, Yuqing Wang, Molei Tao Momentum Stiefel Optimizer, with Applications to Suitably-Orthogonal Attention, and Optimal Transport
ICLR Zhang-Wei Hong, Pulkit Agrawal, Rémi Tachet des Combes, Romain Laroche Harnessing Mixed Offline Reinforcement Learning Datasets via Trajectory Weighting
ICLR Meng Cao, Mehdi Fatemi, Jackie Chi Kit Cheung, Samira Shabanian Systematic Rectification of Language Models via Dead-end Analysis
ICLR Wentao Zhang, Yexin Wang, Zhenbang You, Meng Cao, Ping Huang, Jiulong Shan, Zhi Yang, Bin Cui Information Gain Propagation: a new way to Graph Active Learning with Soft Labels
ICLR Alexandre Perez-Lebel, Marine Le Morvan, Gaël Varoquaux Beyond calibration: estimating the grouping loss of modern neural networks
ICLR Jianwen Xie, Yaxuan Zhu, Jun Li, Ping Li A Tale of Two Flows: Cooperative Learning of Langevin Flow and Normalizing Flow Toward Energy-Based Model
ICLR Amrith Setlur, Don Dennis, Benjamin Eysenbach, Aditi Raghunathan, Chelsea Finn, Virginia Smith, Sergey Levine Bitrate-Constrained DRO: Beyond Worst Case Robustness To Unknown Group Shifts
ICLR Tim Z. Xiao, Robert Bamler Trading Information between Latents in Hierarchical Variational Autoencoders
ICLR Zixuan Ke, Yijia Shao, Haowei Lin, Tatsuya Konishi, Gyuhak Kim, Bing Liu Continual Pre-training of Language Models
ICLR Zhang-Wei Hong, Ge Yang, Pulkit Agrawal Bilinear value networks
ICLR Hanrong Ye, Dan Xu Joint 2D-3D Multi-Task Learning on Cityscapes-3D: 3D Detection, Segmentation, and Depth Estimation
ICLR Mohit Vaishnav, Thomas Serre GAMR: A Guided Attention Model for (visual) Reasoning
ICLR Noam Levi, Itay M. Bloch, Marat Freytsis, Tomer Volansky Noise Injection Node Regularization for Robust Learning
ICLR Paul F. Jaeger, Carsten T. Lüth, Lukas Klein, Till J. Bungert A Call to Reflect on Evaluation Practices for Failure Detection in Image Classification
ICLR Thomas M. Sutter, Laura Manduchi, Alain Ryser, Julia E. Vogt Learning Group Importance using the Differentiable Hypergeometric Distribution
ICLR AmirEhsan Khorashadizadeh, Anadi Chaman, Valentin Debarnot, Ivan Dokmanić FunkNN: Neural Interpolation for Functional Generation
ICML Ramki Gummadi, Saurabh Kumar, Junfeng Wen, Dale Schuurmans A Parametric Class of Approximate Gradient Updates for Policy Optimization
ICML Joshua P. Zitovsky, Daniel de Marchi, Rishabh Agarwal, Michael R. Kosorok Revisiting Bellman Errors for Offline Model Selection
ICML Jiayin Jin, Zeru Zhang, Yang Zhou, Lingfei Wu Input-agnostic Certified Group Fairness via Gaussian Parameter Smoothing
ICML Ching-Yao Chuang, Stefanie Jegelka, David Alvarez-Melis InfoOT: Information Maximizing Optimal Transport
ICML Ilgee Hong, Sen Na, Michael W. Mahoney, Mladen Kolar Constrained Optimization via Exact Augmented Lagrangian and Randomized Iterative Sketching
ICML Matthew Fahrbach, Adel Javanmard, Vahab Mirrokni, Pratik Worah Learning Rate Schedules in the Presence of Distribution Shift
ICML Samuele Marro, Michele Lombardi Computational Asymmetries in Robust Classification
ICML Nicolas Chopin, Andras Fulop, Jeremy Heng, Alexandre H. Thiery Computational Doob’s h-transforms for Online Filtering of Discretely Observed Diffusions
ICML Wentao Zhang, Zeang Sheng, Mingyu Yang, Yang Li, Yu Shen, Zhi Yang, Bin Cui NAFS: A Simple yet Tough-to-beat Baseline for Graph Representation Learning
ICML Disha Shrivastava, Hugo Larochelle, Daniel Tarlow Repository-Level Prompt Generation for Large Language Models of Code
ICML Chenlu Ye, Wei Xiong, Quanquan Gu, Tong Zhang Corruption-Robust Algorithms with Uncertainty Weighting for Nonlinear Contextual Bandits and Markov Decision Processes
ICML Anas Barakat, Ilyas Fatkhullin, Niao He Reinforcement Learning with General Utilities: Simpler Variance Reduction and Large State-Action Space
ICML Alberto Maria Metelli, Francesco Trovò, Matteo Pirola, Marcello Restelli Stochastic Rising Bandits
ICML Idan Shenfeld, Zhang-Wei Hong, Aviv Tamar, Pulkit Agrawal TGRL: An Algorithm for Teacher Guided Reinforcement Learning
ICML Benjamin Dupuis, George Deligiannidis, Umut Şimşekli Generalization Bounds with Data-dependent Fractal Dimensions
ICML Wanrong Zhang, Ruqi Zhang DP-Fast MH: Private, Fast, and Accurate Metropolis-Hastings for Large-Scale Bayesian Inference
ICML Zixuan Ni, Longhui Wei, Siliang Tang, Yueting Zhuang, Qi Tian Continual Vision-Language Representation Learning with Off-Diagonal Information
ICML Manuel Nonnenmacher, Lukas Oldenburg, Ingo Steinwart, David Reeb Utilizing Expert Features for Contrastive Learning of Time-Series Representations
ICML Shih-Yang Liu, Zechun Liu, Kwang-Ting Cheng Oscillation-free Quantization for Low-bit Vision Transformers
ICML Siqi Liu, Marc Lanctot, Luke Marris, Nicolas Heess Simplex Neural Population Learning: Any-Mixture Bayes-Optimality in Symmetric Zero-sum Games
ICML Guanghui Qin, Benjamin Van Durme Nugget: Neural Agglomerative Embeddings of Text
ICML Marc Härkönen, Markus Lange-Hegermann, Bogdan Raiţă Gaussian Process Priors for Systems of Linear Partial Differential Equations with Constant Coefficients
ICML Xiyao Wang, Wichayaporn Wongkamjan, Furong Huang Live in the Moment: Learning Dynamics Model Adapted to Evolving Policy
ICML Tanvir Islam, Peter Washington Personalized Prediction of Recurrent Stress Events Using Self-Supervised Learning on Multimodal Time-Series Data
ICML Krishna Pillutla, Kshitiz Malik, Abdelrahman Mohamed, Michael Rabbat, Maziar Sanjabi, Lin Xiao Federated Learning with Partial Model Personalization
ICML Jaesik Yoon, Yi-Fu Wu, Heechul Bae, Sungjin Ahn An Investigation into Pre-Training Object-Centric Representations for Reinforcement Learning
ICML Mehrdad Ghadiri, Matthew Fahrbach, Gang Fu, Vahab Mirrokni Approximately Optimal Core Shapes for Tensor Decompositions
ICML Arpit Bansal, Ping-yeh Chiang, Michael Curry, Rajiv Jain, Curtis Wigington, Varun Manjunatha, John P Dickerson, Tom Goldstein Certified Neural Network Watermarks with Randomized Smoothing
ICML Mohamad Amin Mohamadi, Wonho Bae, Danica J. Sutherland A Fast, Well-Founded Approximation to the Empirical Neural Tangent Kernel
ICML Chuyang Ke, Jean Honorio Exact Inference in High-order Structured Prediction
ICML Wentao Zhang, Zheyu Lin, Yu Shen, Yang Li, Zhi Yang, Bin Cui DFG-NAS: Deep and Flexible Graph Neural Architecture Search
ICML Tongzhou Wang, Antonio Torralba, Phillip Isola, Amy Zhang Optimal Goal-Reaching Reinforcement Learning via Quasimetric Learning
ICML Yi-Fan Zhang, Xue Wang, Kexin Jin, Kun Yuan, Zhang Zhang, Liang Wang, Rong Jin, Tieniu Tan AdaNPC: Exploring Non-Parametric Classifier for Test-Time Adaptation
ICML Litian Liang, Yaosheng Xu, Stephen McAleer, Dailin Hu, Alexander Ihler, Pieter Abbeel, Roy Fox Reducing Variance in Temporal-Difference Value Estimation via Ensemble of Deep Networks
ICML Mohammed Nowaz Rabbani Chowdhury, Shuai Zhang, Meng Wang, Sijia Liu, Pin-Yu Chen Patch-level Routing in Mixture-of-Experts is Provably Sample-efficient for Convolutional Neural Networks
ICML Gal Leibovich, Guy Jacob, Or Avner, Gal Novik, Aviv Tamar Learning Control by Iterative Inversion
ICML Jiayin Jin, Jiaxiang Ren, Yang Zhou, Lingjuan Lyu, Ji Liu, Dejing Dou Accelerated Federated Learning with Decoupled Adaptive Optimization
ICML Krzysztof Choromanski, Arijit Sehanobish, Han Lin, Yunfan Zhao, Eli Berger, Tetiana Parshakova, Alvin Pan, David Watkins, Tianyi Zhang, Valerii Likhosherstov, Somnath Basu Roy Chowdhury, Avinava Dubey, Deepali Jain, Tamas Sarlos, Snigdha Chaturvedi, Adrian Weller Efficient Graph Field Integrators Meet Point Clouds
ICML Simone Parisi, Aravind Rajeswaran, Senthil Purushwalkam, Abhinav Gupta The Unsurprising Effectiveness of Pre-Trained Vision Models for Control
Table C2: Papers excluded from the analysis.
Conference Paper Title
AAAI Neuro-symbolic Rule Learning in Real-world Classification Tasks
Generalization Bounds for Inductive Matrix Completion in Low-noise Settings
NeurIPS A General Framework for Robust G-Invariance in G-Equivariant Networks
CLIFT: Analysing Natural Distribution Shift on Question Answering Models in Clinical Domain
Partial Counterfactual Identification of Continuous Outcomes with a Curvature Sensitivity Model
Attacks on Online Learners: a Teacher-Student Analysis
Learning Feynman Diagrams using Graph Neural Networks
Function Classes for Identifiable Nonlinear Independent Component Analysis
Blackbox Attacks via Surrogate Ensemble Search
Censored Quantile Regression Neural Networks for Distribution-Free Survival Analysis
Semi-Discrete Normalizing Flows through Differentiable Tessellation
Online Decision Mediation
Exact Generalization Guarantees for (Regularized) Wasserstein Distributionally Robust Models
Deep Learning with Kernels through RKHM and the Perron-Frobenius Operator
Bridging RL Theory and Practice with the Effective Horizon
Reliable learning in challenging environments
ICLR Brain-like representational straightening of natural movies in robust feedforward neural networks
Broken Neural Scaling Laws
Parametrizing Product Shape Manifolds by Composite Networks
Tier Balancing: Towards Dynamic Fairness over Underlying Causal Factors
Guiding continuous operator learning through Physics-based boundary constraints
Scaling Laws For Deep Learning Based Image Reconstruction
Probabilistically Robust Recourse: Navigating the Trade-offs between Costs and Robustness in Algorithmic Recourse
Domain Adaptation via Minimax Entropy for Real/Bogus Classification of Astronomical Alerts
ICML Why Target Networks Stabilise Temporal Difference Methods
Nonlinear Advantage: Trained Networks Might Not Be As Complex as You Think
HyperImpute: Generalized Iterative Imputation with Automatic Model Selection
Note: The papers listed are excluded from the analysis due to tex compilation errors, such as bibtex errors.